john is 8 years older than his sister. 10 years ago the sum of there ages were 82.find the age of each.

Answer by stanbon(75887)

john is 8 years older than his sister. 10 years ago the sum of there ages were 82.find the age of each.
======
Equations:
j = s + 8
j-10 + s-10 = 82
----
Substitute for "j" and solve for "s":
s+8-10 + s - 10 = 82
2s -12 = 82
2s = 94
s = 47 (sister's age now)
----
Solve for "j" using j = s + 8
j = 47+8 = 55 (john's age now)
================================
Cheers,
Stan H.
==============